OpenStudy (anonymous):

Use the slope of 2 and point (-6, -3) to graph the line.

OpenStudy (anonymous):

@AlexandervonHumboldt2 @benlindquist @baru @dayakar @ganeshie8 @iGreen @jigglypuff314 @nincompoop @ParthKohli @phi @Qwertty123 @Rushwr @SamsungFanBoy @TrojanPoem @tkhunny @YoungStudier

OpenStudy (priyar):

\[(y-y1)/(x-x1) = m\] where m is the slope..

OpenStudy (priyar):

the point (x1,y1) has already been given to u

OpenStudy (priyar):

what r u getting @Lelenny?

OpenStudy (anonymous):

So I just need to plug in the numbers?

OpenStudy (anonymous):

(-3 - -3(1)) / (-3 - -3(1))

OpenStudy (trojanpoem):

No , let point (x,y) be on the line use the formula of slope dy/dx = slope slope = 2 point (-6,-3) , (x,y) and find the equation then turn it into y = mx + b (slope-intercept) and graph using x ,y intercept

OpenStudy (anonymous):

Still a little confused.

OpenStudy (trojanpoem):

-3 - y/ -6 - x = 2 3+y/6+x = 2 3+y = 12 + 2x y = 2x + 9 y-intercept : 9 x intercept : -9/2 

OpenStudy (anonymous):

So therefore, it's " https://gyazo.com/2bc582e97ca6ecc196b49849d0787c7b " , correct?